# CodeEnhance: A Codebook-Driven Approach for Low-Light Image Enhancement

### v1

@@ Abstract
Low-light image enhancement (LLIE) aims to improve low-illumination images. However, existing methods face two challenges: (1) uncertainty in restoration from diverse brightness degradations; (2) loss of texture and color information caused by noise suppression and light enhancement. In this paper, we propose a novel enhancement approach, CodeEnhance, by leveraging quantized priors and image refinement to address these challenges. In particular, we reframe LLIE as learning an image-to-code mapping from low-light images to discrete codebook, which has been learned from high-quality images. To enhance this process, a Semantic Embedding Module (SEM) is introduced to integrate semantic information with low-level features, and a Codebook Shift (CS) mechanism, designed to adapt the pre-learned codebook to better suit the distinct characteristics of our low-light dataset. Additionally, we present an Interactive Feature Transformation (IFT) module to refine texture and color information during image reconstruction, allowing for interactive enhancement based on user preferences. Extensive experiments on both real-world and synthetic benchmarks demonstrate that the incorporation of prior knowledge and controllable information transfer significantly enhances LLIE performance in terms of quality and fidelity. The proposed CodeEnhance exhibits superior robustness to various degradations, including uneven illumination, noise, and color distortion.
@@ I Introduction
Classic LLIE methods based on histogram equalization [7] and retinex theories [8][9], are effective in enhancing image brightness. However, these methods neglect image content and color, resulting in a distortion of the output image. Recently, deep learning-based methods [10][11] have been leveraged due to their powerful expressiveness. The popular methods are mainly based on adversarial training [12][13] or image-to-image [14][15] framework, both of which aim to learn the mapping between LL and HQ images. Nevertheless, existing deep learning-based methods encounter limited generalization in real night scenes that contain complex light sources and the intricate illumination. Concretely, (1) The inherent uncertainty in the restoration process stems from addressing diverse brightness degradations, posing a considerable challenge to accurate reconstruction. (2) The loss of texture and color information arises due to the intricate balance required in simultaneous noise suppression and light enhancement, adding complexity to the image enhancement task.
@@
To overcome these challenges, we propose a novel approach named CodeEnhance by feature matching with quantized priors and image refinement. Previous methods usually learn an LL-HQ image mapping, which typically involves a vast parameter space that brings uncertainty for the learning [16]. The key to reducing the uncertainty is to shrink the parameter space. Therefore, we reframe the LLIE task as an image-to-code paradigm by predicting the codebook indices, followed by the pretrained decoder to generate the output. This approach efficiently reduces the parameter space, which alleviates uncertainties in the restoration process and improves robustness against various brightness degradations. Besides, we use the shallow features and introduce reference HQ images to compensate for the texture and color information in the output images, respectively. The rich texture information in shallow features [17] is used to assist the reconstruction. Moreover, the reference images provide valuable guidance for enhancing visual perception in the output.
@@
Specifically, we develop a Semantic Embedding Module (SEM) to integrate semantic information and low-level features in the encoder of CodeEnhance. The SEM plays a crucial role in bridging the semantic gap and ensures effective feature alignment between the encoder and the codebook. Moreover, to address distribution shifts across datasets, we introduce a Codebook Shift (CS) mechanism. This mechanism is designed to adapt the pre-learned codebook from one dataset to better suit the distinct characteristics of our LLIE dataset, ensuring distribution consistency and emphasizing the relevant prior in feature matching. To enhance the restoration of texture and color, we design an Interactive Feature Transformation (IFT) module to fine-tune the texture, color, and brightness of the output image. The IFT consists of two main components: Texture Feature Transformation (TFT) module and Controllable Perceptual Transformation (CPT) module. The TFT module utilizes low-level features from the encoder to refine the details, and the CPT module leverages information from reference images to supplement color information and provide a reference standard for controllable enhancement. By incorporating these modules, we enable a step-by-step refinement process that improves the texture, color, and brightness of the restored image. This design also allows users to adjust the enhancement according to their visual perception, leading to improved customization and user satisfaction.
@@
In summary, CodeEnhance mainly includes two stages. In Stage I, a VQ-GAN [18] is trained using HQ images. In Stage II, we utilize the HQ encoder, the SEM, and the CS to map the LL image to the codebook space. The matched codes are then fed into the frozen HQ decoder and the IFT to generate the enhanced image. Our contributions are listed as follows:
@@
  • Extensive experiments demonstrate that our proposed method achieves state-of-the-art performance on various benchmarks, including LOL, FiveK, and LSRW.
II Related Work
@@ II-A Low-Light Image Enhancement
Images captured in low-light environments often lack important visual details and exhibit poor image quality [19]. These types of images can negatively impact the viewer’s experience and hinder their comprehension of the image content for further analysis. To address this issue, early researchers explored histogram equalization technology [7], which involves adjusting the illumination and contrast of the image by equalizing pixel intensity. They devised various LLIE methods that focused on the overall image perspective [20], cumulative function [21], and penalty terms [22]. In addition, there are many methods based on Retinex theory [9], which decomposes the image into two components: illumination and image reflection [8][23]. With the advancements in deep learning technology, LLNet [6] has successfully integrated it into the field of LLIE for the first time by utilizing stacked autoencoders. To further improve image quality, multi-branch [24], and multi-stage [25] LLIE networks are developed, considering illumination recovery, noise suppression, and color refinement. SCI [26] uses a unique network structure in the training phase, where multiple stages share weights. However, during testing, only one sub-network is used. SNR [2] employs the PSNR distribution map of the image to guide network feature learning and fusion. SMG [3] incorporates image structure information to enhance the output image’s quality. Moreover, by combining Retinex theory with deep learning, URetinexNet [27] formulates the decomposition problem of Retinex as an implicit prior regularization model. Retinexformer [28] uses illumination to guide the Transformer [29] to learn the global illumination information of the image. However, these LLIE methods rely on LL image information for enhancement. This reliance makes them vulnerable to uncertain factors like noise and light sources, causing image artifacts, loss of details, and color distortion. To address these challenges, we propose a novel model that leverages high-quality prior knowledge to enhance its robustness against such uncertain factors. And a new feature transformation module is introduced to enable the algorithm to better handle variations in the input image.
@@
II-B Discrete Codebook Learning
Discrete codebook learning was initially introduced in the VQ-VAE [30]. After that, VQ-GAN [18] incorporates codebook learning into the GAN framework, enabling the generation of high-quality images. In low-level tasks, codebook learning is employed to mitigate uncertainty during the model learning process by transforming the operational space from the image into compact proxy space [16]. To improve feature matching, FeMaSR [31] introduces residual shortcut connections, RIDCP [32] proposes a controllable feature matching operation, and CodeFormer [16] presents a Transformer-based prediction network to obtain the codebook index, respectively. Moreover, LARSR [33] proposes a local autoregressive super-resolution framework based on the learned codebook. VQFR [34] designs parallel decoders to fuse low-level features from the encoder. Building upon the research above, our approach aims to redefine the low-light image enhancement task by introducing the codebook priors and learning the mapping between images and codebook indexes. Additionally, we propose a CS mechanism to fine-tune the original priors. This makes the priors better suitable for different datasets and enables our method to handle various illumination intensities.
III Methodology
The core idea of our CodeEnhance is to improve the mapping from LL images to HQ images by exploiting a discrete representation space while refining the texture and color information for HQ images. The training process of the CodeEnhance involves two stages. In Stage I, a VQ-GAN [18] is trained to learn a discrete codebook prior and its corresponding decoder. As shown in Fig. 2, an SEM is first designed to ensure effective feature alignment between the learnable HQ encoder and the codebook in the Stage II. Then, a CS mechanism is introduced to fine-tune the codebook based on the Stage II dataset to improve feature matching. Finally, we present an IFT module to improve high-quality details and refine visual perceptual.
@@ III-A High-quality Codebook Learning
III-B Feature Matching via Semantic Embedding and Codebook Shift
This section focuses on optimizing feature matching in the codebook by designing a semantic-embedded encoder ${\rm SEE}(\cdot)$ comprising an LL encoder and a key component SEM. Additionally, we propose a Codebook Shift (CS) mechanism ${\rm CS}(\cdot)$ to ensure distribution consistency between the codebook and the current dataset. These techniques improve accuracy and robustness in feature matching. The quantized feature $\mathbf{Z}_{ll}^{q}\in\mathbb{R}^{m\times n\times d}$ of LL image $I_{ll}$ can be obtained by:
$\hat{z}^{(i,j)}_{q}=\mathop{\arg\min}\limits_{c_{k}\in\mathbf{C}}\|{\rm SEE}(I_{ll})^{(i,j)}-{\rm CS}(\mathbf{C})_{k}\|_{2}.$ (3)
@@ III-B Feature Matching via Semantic Embedding and Codebook Shift
As illustrated in Figs. 2 and 3, we use a pre-trained semantic segmentation network (e.g., DeepLab v3 [35]) to extract the semantic feature $\mathbf{F}_{se}$ from the input image. Moreover, we use the learnable HQ encoder to obtain $\mathbf{F}_{ll}$. Subsequently, the $\mathbf{F}_{ll}$ and $\mathbf{F}_{se}$ are sent to the SEM. Within SEM, $\mathbf{F}_{se}$ and $\mathbf{F}_{ll}$ are fused at first, followed by using fused features and $\mathbf{F}_{se}$ to compute an attention-weight map. This map is then applied to the fused features to suppress noise information and obtain semantically embedded features. Lastly, the features are processed by a Multi-Layer Perceptron (MLP) [29]. The SEM can be represented as:
@@ III-C Training Objectives
$\mathcal{L}_{fema}$ $=\beta\|\mathbf{\hat{Z}}_{l}-\mathrm{sg}(\mathbf{Z}_{gt})\|_{2}^{2}+\|\phi(\mathbf{\hat{Z}}_{l})-\phi(\mathrm{sg}(\mathbf{Z}_{gt}))\|_{2}^{2}$
$+\|\mathbf{Z}_{ll}^{q}-\mathrm{sg}(\mathbf{Z}_{gt}^{q})\|_{2}^{2},$ (9)
where $\phi(\cdot)$ is used to calculate the gram matrix of features. $\mathbf{\hat{Z}}_{l}$ and $\mathbf{Z}_{gt}$ represent the latent features of LL images and ground truth, respectively. $\mathbf{Z}_{ll}^{q}$ and $\mathbf{Z}_{gt}^{q}$ denote the quantized features of LL images and ground truth, respectively.
@@ IV Experiments
Methods Metrics LIME[42] JED[43] RetinexNet[1] KinD[44] Zero[45] EnGAN[12] SNR[2] MIRNetv2[4] PairLIE[46] SMG[3] Ours
@@
IV-A Implementation Details
IV-B Datasets and Evaluation Metrics
Low-light Datasets. We use LSRW [38], FiveK [41], LOL [1], and SynLL [24] datasets to evaluate the proposed method. The LSRW contains Huawei and Nikon datasets. The Huawei consists of training and testing sets with 2450 and 30, respectively. The Nikon includes 3,150 and 20 image pairs for training and testing, respectively. FiveK [41] dataset consists of 4500 image pairs for training and 500 image pairs for testing. The LOL dataset includes 485 and 15 image pairs for training and testing, respectively. The SynLL [24] benchmark synthesizes 23,431 short/long exposure image pairs by imposing the degradation of illumination, color, and noise. 22,472 and 959 image pairs are used for training and testing.
Evaluation Metrics. We use the most common peak signal-to-noise ratio (PSNR), structural similarity [48] (SSIM), mean abstract error (MAE), and learned perceptual image patch similarity (LPIPS) [40] to measure the quality of the enhanced image. Unlike PSNR and MAE, LPIPS takes into account the human visual system’s perception of similarity, which provides a more accurate assessment of the perceived similarity between images.
@@ IV-B Datasets and Evaluation Metrics
Figure 10: Visual results on the real scene. The results demonstrate the superior performance of our proposed CodeEnhance method. Particularly, it excels in preserving texture details and effectively handling gradient change regions.

### v3

@@ Abstract
Low-light image enhancement (LLIE) aims to improve low-illumination images. However, existing methods face two challenges: (1) uncertainty in restoration from diverse brightness degradations; (2) loss of texture and color information caused by noise suppression and light enhancement. In this paper, we propose a novel enhancement approach, CodeEnhance, by leveraging discrete codebook priors and image refinement to address these challenges. In particular, we reframe LLIE as learning an image-to-code mapping from low-light images to discrete codebook, which has been learned from high-quality images. To enhance this process, a Semantic Embedding Module (SEM) is introduced to integrate semantic information with low-level features, and a Codebook Shift (CS) mechanism, designed to adapt the pre-learned codebook to better suit the distinct characteristics of our low-light dataset. Additionally, we present an Interactive Feature Transformation (IFT) module to refine texture and color information during image reconstruction, allowing for interactive enhancement based on user preferences. Extensive experiments on both real-world and synthetic benchmarks demonstrate that the incorporation of prior knowledge and controllable information transfer significantly enhances LLIE performance in terms of quality and fidelity. The proposed CodeEnhance exhibits superior robustness to various degradations, including uneven illumination, noise, and color distortion. Our project page is https://github.com/csxuwu/CodeEnhance.
@@ I Introduction
Classic LLIE methods based on histogram equalization [7] and retinex theories [8][9], are effective in enhancing image brightness. However, these methods neglect image content and color, resulting in a distortion of the output image [10][11]. Recently, deep learning-based methods have been leveraged due to their powerful expressiveness. The popular methods are mainly based on adversarial training [12] or image-to-image [13][14] framework, both of which aim to learn the mapping between LL and HQ images. Nevertheless, existing deep learning-based methods encounter limited generalization in real night scenes that contain complex light sources and intricate illumination. Concretely, (1) Uncertainty in LLIE: Real-world nighttime scenes exhibit a broad spectrum of brightness degradations, making it inherently challenging to achieve accurate reconstruction. This diversity introduces uncertainty into the restoration process, as conventional methods often cannot capture the complexity of varying illumination conditions. (2) Tradeoffs in LLIE: In low-light scenarios, removing noise often comes at the expense of texture and color information. As shown in Fig. 1, current methods frequently over-smooth or distort the enhanced image, underscoring the challenge of balancing noise suppression with the need to preserve fine details and natural color.
@@
To overcome these challenges, we propose a novel approach named CodeEnhance by feature matching with quantized priors and image refinement. Previous methods usually learn an LL-HQ image mapping, which typically involves a vast parameter space that introduces uncertainty for the learning [15]. The key to reducing the uncertainty is to shrink the parameter space. Therefore, we reframe the LLIE task as an image-to-code paradigm by extracting image features and converting them to the codebook indices, followed by the pretrained decoder to generate the output. This approach efficiently reduces the parameter space, which alleviates uncertainties in the restoration process and improves robustness against various brightness degradations. Besides, we use the shallow features to compensate for texture information in the reconstruction process. We further introduce HQ images to guide accurate color enhancement in the output, thereby enhancing visual perception.
@@
Specifically, a Semantic Embedding Module (SEM) is developed to integrate semantic information and low-level features in the encoder of CodeEnhance. The SEM is crucial for bridging the semantic gap between the encoder output and the codebook, leading to effective feature matching. Moreover, to address distribution shifts across datasets, we introduce a Codebook Shift (CS) mechanism. This mechanism is designed to adapt the pre-learned codebook from one dataset to better suit the distinct characteristics of our LLIE dataset. It ensures distribution consistency and emphasizes relevant priors in feature matching. To improve the restoration of texture and color, we design an Interactive Feature Transformation (IFT) module to fine-tune the texture, color, and brightness of the output image. The IFT consists of two main components: Texture Feature Transformation (TFT) module and Controllable Perceptual Transformation (CPT) module. The TFT module utilizes low-level features from the encoder to refine the details, and the CPT module leverages information from reference images to supplement color information and provide a reference standard for controllable enhancement. By incorporating these modules, we enable a step-by-step refinement process that improves the texture, color, and brightness of the restored image. This design also allows users to adjust the enhancement according to their visual perception, leading to improved customization and user satisfaction.
@@
In summary, CodeEnhance mainly includes two stages. In Stage I, a VQ-GAN [16] is trained in HQ images by self-reconstruction. In Stage II, we utilize the HQ encoder, the SEM, and the CS to map the LL image to the codebook space. The matched codes are then fed into the frozen HQ decoder and the IFT to generate the enhanced image. Our contributions are listed as follows:
@@
  • Extensive experiments demonstrate that our proposed method achieves state-of-the-art performance on various benchmarks, including LOL, FiveK, LSRW, LIME, MEF, and NPE.
The remainder of this paper is organized as follows. Section II reviews the related works. Section III presents the model design in detail. Section IV conducts experiments and analyzes the results. Finally, we conclude the paper in Section V.
II Related Works
@@ II-A Low-Light Image Enhancement
Images captured in low-light environments often lack important visual details and exhibit poor image quality [17][18]. These types of images can negatively impact the viewer’s experience and hinder their comprehension of the image content for further analysis. To address this issue, early researchers explored histogram equalization technology [7], which involves adjusting the illumination and contrast of the image by equalizing pixel intensity. They devised various LLIE methods that focused on the overall image perspective [19], cumulative function [20], and penalty terms [21]. In addition, many methods based on Retinex theory [9] decompose the image into two components: illumination and image reflection [8]. With the advancements in deep learning technology [22][23][24], LLNet [2] has successfully integrated it into the field of LLIE for the first time utilizing stacked autoencoders. To further improve image quality, multi-branch [11][25], and multi-stage [26] LLIE networks are developed, considering illumination recovery, noise suppression, and color refinement. SCI [27] uses a unique network structure in the training phase, where multiple stages share weights. However, during testing, only one subnetwork is used. SNR [4] employs the PSNR distribution map of the image to guide network feature learning and fusion. SMG [5] incorporates information about the image structure to enhance the quality of the output image. Moreover, by combining Retinex theory with deep learning, URetinexNet [28] and DA-DRN [29] formulated the decomposition problem of Retinex as an implicit prior regularization model. Retinexformer [30] learns the global illumination information by using illumination-guided Transformer [31]. END [32] enhances low-light images by dual prior guidance. With the development of diffusion model [33][34], GSAD [35], QuadPrior[36] and JoRes [37] leverage the diffusion model to perform LLIE.
However, these LLIE methods rely on LL image information for enhancement. This reliance makes them vulnerable to uncertain factors like noise and light sources, causing image artifacts, loss of details, and color distortion. To address these challenges, we propose a novel model that leverages high-quality prior knowledge to enhance its robustness against such uncertain factors. And a new feature transformation module is introduced to enable the algorithm to better handle variations in the input image.
II-B Discrete Codebook Learning
Discrete codebook learning was initially introduced in the VQ-VAE [38]. After that, VQ-GAN [16] incorporates codebook learning into the GAN framework, enabling the generation of high-quality images. In low-level tasks, codebook learning is employed to mitigate uncertainty during the model learning process by transforming the operational space from the image into compact proxy space [15]. To improve feature matching, FeMaSR [39] introduces residual shortcut connections, RIDCP [40] proposes a controllable feature matching operation, and CodeFormer [15] presents a Transformer-based prediction network to obtain the codebook index, respectively. Moreover, LARSR [41] proposes a local autoregressive super-resolution framework based on the learned codebook. VQFR [42] designs parallel decoders to fuse low-level features from the encoder. CodeBGT [43] introduces the codebook to improve LLIE model performance.
Building upon the research above, our approach aims to redefine the low-light image enhancement task by introducing the codebook priors and learning the mapping between images and codebook indexes. Additionally, we propose a CS mechanism to fine-tune the original priors. This makes the priors better suitable for different datasets and enables our method to handle various illumination intensities.
III Methodology
The core idea of the proposed CodeEnhance is to minimize restoration uncertainty by reframing the LLIE as an image-to-code paradigm. The training process for CodeEnhance involves two stages. In Stage I, a VQ-GAN [16] is trained to learn a discrete codebook prior and its corresponding decoder. As shown in Fig. 2, an SEM is first designed to ensure effective feature alignment between the learnable HQ encoder and the codebook in the Stage II. Then, a CS mechanism is introduced to fine-tune the codebook based on the Stage II dataset to improve feature matching. Finally, we present an IFT module to improve high-quality details and refine visual perceptual.
@@ III-A High-quality Codebook Learning
III-B Feature Matching via Semantic Embedding and Codebook Shift
This section focuses on optimizing feature matching in the codebook by designing a semantic-embedded encoder ${\rm SEE}(\cdot)$ comprising an HQ encoder ${\rm E}(\cdot)$ and a key component Semantic Embedding Module (SEM) ${\rm SEM}(\cdot)$. Additionally, we propose a Codebook Shift (CS) mechanism ${\rm CS}(\cdot)$ to ensure distribution consistency between the codebook and the current dataset. These techniques improve accuracy and robustness in feature matching. The Formula 1 can be redefined as follows:
$z^{(i,j)}_{q}=\mathop{\arg\min}\limits_{c_{k}\in\mathbf{C}}\|{\rm SEE}(I_{ll})^{(i,j)}-{\rm CS}(c_{k})\|_{2},$ (3)
${\rm SEE}(I_{ll})={\rm SEM}({\rm E}(I_{ll}),{\rm DL}(I_{ll}))$
where ${\rm DL}(\cdot)$ represents a pre-trained DeepLab v3+ [44] network, which is used to extract semantic information.
@@ III-B Feature Matching via Semantic Embedding and Codebook Shift
As illustrated in Figs. 2 and 3, we use a pre-trained semantic segmentation network (e.g., DeepLab v3+ [44]) to extract the semantic feature $\mathbf{F}_{se}$ from the input image. Moreover, we use the learnable HQ encoder to obtain $\mathbf{F}_{ll}$. Subsequently, the $\mathbf{F}_{ll}$ and $\mathbf{F}_{se}$ are sent to the SEM. Within SEM, $\mathbf{F}_{se}$ and $\mathbf{F}_{ll}$ are fused at first, followed by using fused features and $\mathbf{F}_{se}$ to compute an attention-weight map. This map is then applied to the fused features to suppress noise information and obtain semantically embedded features. Lastly, the features are processed by a Multi-Layer Perceptron (MLP) [31]. The SEM can be represented as:
@@
III-C Image Refinement via Interactive Feature Transformation
We develop Interactive Feature Transformation (IFT) that balances recovering texture details, contrast, and brightness. As shown in Fig. 2, the IFT allows us to control the information transmitted to the decoder by Texture Feature Transformation (TFT) and Controllable Perceptual Transformation (CPT). The IFT can be formulated as follows:
$\mathbf{F}_{IFT}=\mathbf{F}_{TFT}+\mathbf{F}_{CPT},$ (6)
where $\mathbf{F}_{TFT}$ and $\mathbf{F}_{CPT}$ are the output of the TFT and CPT, respectively.
Texture Feature Transformation. The low-level feature $\mathbf{F}_{e}$ from the HQ encoder contains texture information [48]. Besides, noise in the low-level feature may affect the quality of reconstruction. To address this problem, we design a TFT to effectively incorporate significant texture information and compensate for the loss of details. As shown in Fig. 2, it first fuses $\mathbf{F}_{e}$ and the decoder features $\mathbf{F}_{d}$, and then obtains the affine transformation parameters $\bm{\alpha}$ and $\bm{\beta}$ to mitigate the impact of noise and extract texture. Thereby, the TFT can adaptively refine the decoder features by feature affine transformation [49] as follows:
$\mathbf{F}_{TFT}$ $=\bm{\alpha}\odot\mathbf{F}_{d}+\bm{\beta},$ (7)
$\bm{\alpha},\bm{\beta}$ $=\mathcal{C}([\mathbf{F}_{d},\mathbf{F}_{e}]),$
where $\mathcal{C}(\cdot)$ denotes convolution operation, and $\odot$ is element-wise multiplication.
@@ III-D Training Objectives
$\mathcal{L}_{fema}$ $=\beta\|\mathbf{\hat{Z}}_{ll}-\mathrm{sg}(\mathbf{Z}_{gt}^{q})\|_{2}^{2}+\|\phi(\mathbf{\hat{Z}}_{ll})-\phi(\mathrm{sg}(\mathbf{Z}_{gt}))\|_{2}^{2}$
where $\phi(\cdot)$ is used to calculate the gram matrix of features. $\mathbf{\hat{Z}}_{ll}$ and $\mathbf{Z}_{gt}^{q}$ represent the latent features of LL images and quantized features of ground truth, respectively.
@@
Overall, the proposed CodeEnhance is constructed through two key modules: feature matching via semantic embedding and codebook shift, and image refinement via interactive feature transformation. This approach redefines LLIE as a task of transforming low-light images into a codebook. Next, we will demonstrate the effectiveness of CodeEnhance through extensive experimental validation.
@@ IV Experiments
IV-A Implementation Details
IV-B Datasets and Evaluation Metrics
Low-light Datasets. We use LSRW [47], FiveK [53], LOL [3], and SynLL [11] datasets to evaluate the proposed method. The LSRW contains Huawei and Nikon datasets. The Huawei consists of training and testing sets with 2450 and 30, respectively. The Nikon includes 3,150 and 20 image pairs for training and testing, respectively. FiveK [53] dataset consists of 4500 image pairs for training and 500 image pairs for testing. The LOL dataset includes 485 and 15 image pairs for training and testing, respectively. The SynLL [11] benchmark synthesizes 23,431 short/long exposure image pairs by imposing the degradation of illumination, color, and noise. 22,472 and 959 image pairs are used for training and testing. Additionally, we use unpaired datasets including LIME [54], MEF [10], and NPE [55] to evaluate generalization performance.
Evaluation Metrics. We use the most common Peak Signal-to-Noise Ratio (PSNR), structural similarity [56] (SSIM), Mean Abstract Error (MAE), and Learned Perceptual Image Patch Similarity (LPIPS) [51] to measure the quality of the enhanced image. LPIPS considers the human visual system’s perception of similarity, which provides a more accurate assessment of the perceived similarity between images. In addition, natural Image Quality Evaluator (NIQE) [57] is used to measure results without ground truth.
@@ IV-B Datasets and Evaluation Metrics
Methods Metrics LIME JED RetinexNet KinD Zero EnGAN SNR MIRNetv2 PairLIE SMG QuadPrior Ours
@@
((a)) PSNR $\uparrow$
((b)) SSIM $\uparrow$
((c)) LPIPS $\downarrow$
((d)) MAE $\downarrow$
Figure 6: Panels (a)-(d) correspond to Table I, showing PSNR, SSIM, LPIPS, and MAE across test sets. $\uparrow$ means higher is better, $\downarrow$ means lower is better.
@@
((a)) LIME $\downarrow$
((b)) MEF $\downarrow$
((c)) NPE $\downarrow$
((d)) ExDark $\uparrow$
Figure 11: Panels (a)-(c) correspond to Table II, which show NIQE comparisons across LIME [54], MEF [10], and NPE [55]. Panel (d) corresponds to Table IV, showing mAP on ExDark [1]. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better.
TABLE II: Comparisons on LIME [54], MEF [10], and NPE [55] dataset in terms of NIQE, where the lower the better.
Methods KinD EnlightGAN Zero SNR PairLIE MIRNetv2 SMG CodeBGT QuadPrior Ours
LIME 6.71 3.59 3.78 4.88 4.31 4.12 6.47 4.20 4.58 3.41
MEF 3.17 3.11 3.31 3.47 3.92 3.33 6.18 3.85 4.36 2.95
NPE 3.28 4.36 3.47 4.19 3.67 3.66 5.89 3.52 3.65 3.11